OpenStudy (anonymous):

What is square root of 7 divided by the 4th root of 7? (Like a 4 in front of the square root sign. I only know square root and cubed root)

OpenStudy (abb0t):

Can you write it, there is a little button on the bottom of this little box that says "∑ Equation", click that to write it out.

OpenStudy (anonymous):

Yeah sure it's: \[\sqrt{7}\div \sqrt[4]{7}\]

OpenStudy (triciaal):

when you divide number with the same base subtract the exponents

OpenStudy (triciaal):

remember sq rt = power 1/2 and 4th root = power 1/4 for this problem your base is 7

OpenStudy (anonymous):

So you would subtract 1/2 from 1/4? Sorry, I know this seems simple but this is one of the main things I didn't get in Alg 2

OpenStudy (triciaal):